OpenStudy (anonymous):

Can someone help me with a step by step explanation for this problem? Ann and Tom want to establish a fund for their grandson’s college education. What lump sum must they deposit at a 9.7% annual interest rate, compounded quarterly, in order to have $30,000 in the end of 15 years?

OpenStudy (anonymous):

wishful thinking solve \[30,000=P\left(1+\frac{.097}{12}\right)^{12\times 15}\]for \(P\)

OpenStudy (anonymous):

nope, that was monthly sorry

OpenStudy (anonymous):

solve \[30,000=P\left(1+\frac{.097}{4}\right)^{4\times 15}\]

OpenStudy (anonymous):

When I plugged the equation into my calculator I ended up with 4.210. I'm confused about what I do next. For example, what must I do with the 30000 part? Am I supposed to subtract something? Or am I just over thinking this, and that is indeed the answer?